OpenStudy (anonymous):

Circle A has a diameter of 9 inches, a circumference of 28.26 inches, and an area of 63.585 square inches. The diameter of circle B is 5 inches, the circumference is 15.70 inches, and the area is 19.625 square inches.. Part A: Using the formula for circumference, solve for the value of pi for each circle. (4 points) Part B: Use the formula for area and solve for the value of pi for each circle. (4 points) Part C: What observation can you make about the value of pi for circles A and B? (2 points)

Miracrown (miracrown):

Have you attempted any yet?

OpenStudy (anonymous):

not really. i dont really know about this

OpenStudy (anonymous):

@Miracrown

Miracrown (miracrown):

Let's focus on circle A. With this info, can you find the value of pi being used for circle A?

OpenStudy (anonymous):

3.14

OpenStudy (anonymous):

right?

Miracrown (miracrown):

Circle A: R = 9/2 inches, C = 28.26, A = 63,585

Miracrown (miracrown):

Yes, you're right!

OpenStudy (anonymous):

ok so then next.... ?

Miracrown (miracrown):

Let's now look at circle B: R = 5/2, C = 15.70, A = 19.625

OpenStudy (anonymous):

Pi is the same

OpenStudy (anonymous):

3.14?

Miracrown (miracrown):

Correct :)

OpenStudy (anonymous):

Wow

Miracrown (miracrown):

So it appears the person working these calculations for circles A and B is using the same consistent value for pi or 3.14

Miracrown (miracrown):

See how easy it all was? :P

OpenStudy (anonymous):

Yes, true

OpenStudy (anonymous):

Wow, you're right :)

OpenStudy (anonymous):

But before you leave, can you tell me what i'm supposed to put for them? in one post?

OpenStudy (anonymous):

please?

OpenStudy (anonymous):

I have to show my work :( but im not surehow to do that

Miracrown (miracrown):

Oh, we just solved it. You just simply put it all together, like this: A = 3.14 B = 3.14 C =The person working these calculations for circles A and B is using the same consistent value for pi or 3.14

OpenStudy (anonymous):

For real??? thx :)

OpenStudy (anonymous):

ur the best!!!

OpenStudy (anonymous):

can you help me with some more?

OpenStudy (anonymous):

ill put it in a new post

Miracrown (miracrown):

Sure